# Cropping outperforms dropout as an augmentation strategy for training self-supervised text embeddings

### v1

@@ Appendix B Supplementary tables and figures
SBERT sentence-transformers/all-mpnet-base-v2 (Reimers & Gurevych, 2019) 2021
@@
Table S2: Effect of post-processing transformations. kkNN accuracy on the ICLR dataset using different post-processing transformations of the MPNet mean pooling representation, obtained via the Euclidean and the cosine metrics for finding nearest neighbors, before and after fine-tuning the model on the ICLR dataset. Whitening MPNet’s representation increased the performance before fine-tuning when using the cosine metric for NN search, but it did not when using Euclidean metric. Whitening did not bring any further improvement to the representation after fine-tuning, with any of the metrics.
@@
Table S4: Dataset statistics. Statistics of the datasets used in the experiments of Table 2. The arXiv, bioRxiv, medRxiv, Reddit, and StackExchange datasets are from the P2P clustering tasks of the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2023), and the ICLR dataset is taken from González-Márquez & Kobak (2024). Length refers to the number of characters in each sample text. For the arXiv dataset, we used secondary paper categories (e.g., “cs.AI”) as labels.
@@
Figure S1: Hyperparameter tuning using the ICLR dataset for self-supervised training and kkNN evaluation. (a) Temperature $\tau$ used to scale the similarities in the loss function. (b) Number of consecutive sentences $t$ used in the cropping augmentation. The minibatch size $b$ was adapted depending on $t$ to make it fit into our GPU memory: we used $b=128$ for $t=1$; $b=64$ for $t=2,3,4$; $b=32$ for $t=5,6,7,8,9$; and $b=16$ for $t=10$. (c) Fraction of masked tokens used in addition of the cropping augmentation. (d) Learning rate $\eta$ used by the Adam optimizer.

### v2

@@ A.2 Description of MTEB tasks
Classification
Each of the used datasets consists of a set of text and labels, which can be either topic labels or sentiment labels. The texts from the train and test sets are embedded with the model. The train set embeddings are used to train a logistic regression classifier with 100 maximum iterations, which is scored on the test set. The main metric is linear classification accuracy.
@@ Appendix B Supplementary tables and figures
Table S1: Different base models. This is an extension of Table 1. Different base models (MPNet, BERT, RoBERTa) before and after fine-tuning with our main augmentations. All values in percent, higher is better. Models in columns 2–3, 5–6, and 8–9 were fine-tuned on the ICLR dataset.
(1) (2) (3) (4) (5) (6) (7) (8) (9)
Model MPNet MPNet MPNet BERT BERT BERT RoBERTa RoBERTa RoBERTa
Augmentations Dropout Crops Dropout Crops Dropout Crops
Clustering 27.4 34.7 37.7 32.3 35.1 36.4 21.9 35.4 36.8
Reranking 41.8 46.5 51.9 46.7 47.4 51.5 40.3 47.1 48.8
Retrieval 11.8 24.2 31.8 15.6 21.3 29.4 7.2 24.0 25.0
STS 52.0 65.9 73.4 57.1 59.4 69.6 59.3 69.8 70.2
Classification 61.6 63.3 64.3 65.5 62.6 64.2 63.3 63.5 64.7
Block average 38.9 46.9 51.8 43.4 45.2 50.2 38.4 48.0 49.1
@@
SBERT sentence-transformers/all-mpnet-base-v2 (Reimers and Gurevych, 2019) 2021
@@
Table S3: Effect of post-processing transformations. kkNN accuracy on the ICLR dataset using different post-processing transformations of the MPNet mean pooling representation, obtained via the Euclidean and the cosine metrics for finding nearest neighbors, before and after fine-tuning the model on the ICLR dataset.
@@
Table S5: Dataset statistics. Statistics of the datasets used in the experiments of Table 2. The arXiv, bioRxiv, medRxiv, Reddit, and StackExchange datasets are from the P2P clustering tasks of the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2023), and the ICLR dataset is taken from González-Márquez and Kobak (2024). Length refers to the number of characters in each sample text. For the arXiv dataset, we used secondary paper categories (e.g. “cs.AI”) as labels.
@@
Figure S1: Hyperparameter tuning. kkNN accuracies on the ICLR dataset used for self-supervised training and evaluation, as a function of different hyperparameter values. (a) Temperature $\tau$ used to scale the similarities in the loss function. (b) Number of consecutive sentences $t$ used in the cropping augmentation. The minibatch size $b$ was adapted depending on $t$ to make it fit into our GPU memory: we used $b=128$ for $t=1$; $b=64$ for $t=2,3,4$; $b=32$ for $t=5,6,7,8,9$; and $b=16$ for $t=10$. (c) Fraction of masked tokens used in addition of the cropping augmentation. (d) Learning rate $\eta$ used by the Adam optimizer.
@@
Figure S3: MTEB score across layers. MTEB score using the representation after each layer for MPNet after cropping-based fine-tuning (as in Figure 4c), split by task modality. Layer 0 corresponds to the embedding layer. Asterisks highlight the highest performance in each task modality.